# Simple and Provable Scaling Laws for the Test-Time Compute of Large Language Models

### v4

@@ Appendix B Examples for understanding and comparing the assumptions
Example 2.
We demonstrate a scenario where both Assumptions 1 and 2 hold, but the parameter $\Delta$ in the latter can be much smaller than $p_{\text{comp}}-0.5$ in the former. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=p_{B}=\alpha$, $p_{C}=1-2\alpha$, ${\mathbb{P}}(A\succ B)=p_{\text{comp}}>0.5$, and ${\mathbb{P}}(A\succ C)={\mathbb{P}}(B\succ C)=0.9$. Then we can calculate the average win rate of each candidate as follows:
@@
Example 3.
We demonstrate a scenario where Assumption 1 holds true but Assumption 2 does not. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=0.2,p_{B}=0.2,p_{C}=0.6$, ${\mathbb{P}}(A\succ B)={\mathbb{P}}(A\succ C)=0.6$, and ${\mathbb{P}}(B\succ C)=0.9$, which satisfies Assumption 1. Then we have
@@
In other words, the average win rate of the only correct solution $A$ is lower than that of an incorrect solution $B$, which violates Assumption 2.
Example 4.
We demonstrate a scenario where Assumption 2 holds true but Assumption 1 does not. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=0.2,p_{B}=0.2,p_{C}=0.6$, ${\mathbb{P}}(A\succ B)=0.4,{\mathbb{P}}(A\succ C)=0.9$, and ${\mathbb{P}}(B\succ C)=0.5$, which violates Assumption 1. However, we have
@@
which satisfies Assumption 2 since the only correct solution $A$ has the highest average win rate.
Appendix C Supplementary materials for experiments
C.1 Additional implementation details
@@
During the generation stage, we ask the LLM (via zero-shot CoT prompting [18]) to generate a reasoning process first and then its final answer. For each pairwise comparison during the aggregation stage, we also leverage zero-shot CoT prompting and ask the LLM to think step by step before deciding which solution in the pair is more plausible, unless specified otherwise. Tables C.3 and C.3 at the end of this section include the prompt templates used in our experiments for GPQA and MMLU-Pro-S, both of which are multiple-choice datasets. The prompt templates for MATH-500 are largely the same, only slightly adjusted to account for the desired output formats. Some parts of our prompts, as well as code for parsing LLMs’ responses and extracting the answers for evaluation, are modified from those in the official GitHub repository of MMLU-Pro [^4] https://github.com/TIGER-AI-Lab/MMLU-Pro/tree/main.
@@
C.2 Additional results for the knockout-style algorithm
@@
Figure 6: Accuracy versus the number of initial candidates $N$ for the knockout-style algorithm on MMLU-Pro-S (left) and MATH-500 (right).
@@
Figure 7: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the empirical results for the knockout-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{p}_{\text{comp}}>0.5$, $\widehat{p}_{\text{comp}}\leq 0.5$, $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{gen}}=1$. To the right of each plot is a histogram for $\widehat{p}_{\text{comp}}$. Each problem is represented by a circle if it is solved correctly by the knockout-style algorithm with $N=64$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{gen}}=0$ or $1$, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{p}_{\text{comp}}$ for such problems.
@@
C.3 Additional results for the league-style algorithm
@@
Figure 14: Empirical results for the league-style algorithm on MMLU-Pro-S. Left: accuracy versus the number of initial candidates $N$ for the league-style (LG, solid lines) and knockout-style (KO, dotted lines) algorithms, given the same initial candidates. Right: accuracy versus $M$, the number of subsampled opponents for each candidate, for the league-style algorithm with $N=16$.
@@
Figure 15: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$ that are estimated with the empirical results for the league-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{\Delta}>0$, $\widehat{\Delta}\leq 0$, $\widehat{p}_{\text{cs}}=0$ or $\widehat{p}_{\text{cs}}=1$. To the right of each plot is a histogram for $\widehat{\Delta}$. Each problem is represented by a circle if it is solved correctly by the league-style algorithm with $N=16$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{cs}}=0$ or $1$, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{\Delta}$ for such problems.
@@
Figure 16: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{comp}}$ from the knockout-style algorithm and $\widehat{\Delta}$ from the league-style algorithm (both with $N=16$) using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. The following abbreviations are used for the legend: K — knockout, L — league, R — right, W — wrong. For example, “KW-LR (# = 66)” means that there are 66 problems for which the knockout-style algorithm did wrong while the league-style algorithm did right.
@@
Please read the following multiple-choice questions and provide the most likely correct answer based on the options given. # Question {question} # Output Format “‘ <reason>your step-by-step reasoning proecss</reason> <answer>"the answer is (X)" where X is the correct letter choice</answer> “‘
Table 2: The adopted prompt template for pairwise comparison.
You are an impartial Judge. Given a question and two candidate solutions, your task is to choose which solution answer the question better. Your judgment should be unbiased, without favoring either Solution 1 or 2. —- QUESTION —- {question} —- Solution 1 —- {candidate_a} —- Solution 2 —- {candidate_b} —- OUTPUT FORMAT —- “‘ <compare>compare both candidate solutions step-by-step thoroughly, and double check if there are mistakes in either solution</compare> <winner>Solution 1 or Solution 2 or Tie</winner> “‘

### v5

@@ Appendix C Examples for understanding and comparing the assumptions
Example C.2.
We demonstrate a scenario where both Assumptions 2.1 and 3.1 hold, but the parameter $\Delta$ in the latter can be much smaller than $p_{\text{comp}}-0.5$ in the former. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=p_{B}=\alpha$, $p_{C}=1-2\alpha$, ${\mathbb{P}}(A\succ B)=p_{\text{comp}}>0.5$, and ${\mathbb{P}}(A\succ C)={\mathbb{P}}(B\succ C)=0.9$. Then we can calculate the average win rate of each candidate as follows:
@@
Example C.3.
We demonstrate a scenario where Assumption 2.1 holds true but Assumption 3.1 does not. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=0.2,p_{B}=0.2,p_{C}=0.6$, ${\mathbb{P}}(A\succ B)={\mathbb{P}}(A\succ C)=0.6$, and ${\mathbb{P}}(B\succ C)=0.9$, which satisfies Assumption 2.1. Then we have
@@
In other words, the average win rate of the only correct solution $A$ is lower than that of an incorrect solution $B$, which violates Assumption 3.1.
Example C.4.
We demonstrate a scenario where Assumption 3.1 holds true but Assumption 2.1 does not. Suppose that $\mathcal{Y}=\{A,B,C\}$, where only $A$ is correct. In addition, $p_{A}=0.2,p_{B}=0.2,p_{C}=0.6$, ${\mathbb{P}}(A\succ B)=0.4,{\mathbb{P}}(A\succ C)=0.9$, and ${\mathbb{P}}(B\succ C)=0.5$, which violates Assumption 2.1. However, we have
@@
which satisfies Assumption 3.1 since the only correct solution AA has the highest average win rate.
Appendix D Supplementary materials for experiments
D.1 Additional implementation details
@@
During the generation stage, we ask the LLM (via zero-shot CoT prompting [18]) to generate a reasoning process first and then its final answer. For each pairwise comparison during the aggregation stage, we also leverage zero-shot CoT prompting and ask the LLM to think step by step before deciding which solution in the pair is more plausible, unless specified otherwise. Tables D.3 and LABEL:tab:prompt_comparison at the end of this section include the prompt templates used in our experiments for GPQA and MMLU-Pro-S, both of which are multiple-choice datasets. The prompt templates for MATH-500 are largely the same, only slightly adjusted to account for the desired output formats. Some parts of our prompts, as well as code for parsing LLMs’ responses and extracting the answers for evaluation, are modified from those in the official GitHub repository of MMLU-Pro [^5] https://github.com/TIGER-AI-Lab/MMLU-Pro/tree/main.
@@
Due to the high computational or monetary costs of the experiments, we have run the knockout/league-style algorithm only once for each <model, dataset> combination. To enhance the stability and reliability of the plots in this paper, we take the following approaches:
  • For the knockout-style algorithm, we take advantage of its binary tree structure (shown in Figure 1). After running the algorithm once with $N=64$, we automatically get the results of 64 independent trials for $N=1$, 32 trials for $N=2$, 16 trials for $N=4$, and so on. We have thus taken the average of accuracy values from multiple independent trials for each datapoint (except for the rightmost one) in each scaling curve.
  • For the league-style algorithm, after running it once with $N=16$, we are able to obtain the results of multiple trials for $N=8$ (or any value smaller than 16), each corresponding to 8 randomly sampled candidate solutions and the comparison results among them. Each datapoint (except for the rightmost one) in each scaling curve has been calculated by an average of multiple results obtained this way.
D.2 Additional results for the knockout-style algorithm
@@
Figure 6: Accuracy versus the number of initial candidates NN for the knockout-style algorithm on MMLU-Pro-S (left) and MATH-500 (right).
@@
Figure 7: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the empirical results for the knockout-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{p}_{\text{comp}}>0.5$, $\widehat{p}_{\text{comp}}\leq 0.5$, $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{gen}}=1$. To the right of each plot is a histogram for $\widehat{p}_{\text{comp}}$. Each problem is represented by a circle if it is solved correctly by the knockout-style algorithm with $N=64$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{gen}}=0$ or 11, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{p}_{\text{comp}}$ for such problems.
@@
D.3 Additional results for the league-style algorithm
@@
Figure 14: Empirical results for the league-style algorithm on MMLU-Pro-S. Left: accuracy versus the number of initial candidates NN for the league-style (LG, solid lines) and knockout-style (KO, dotted lines) algorithms, given the same initial candidates. Right: accuracy versus $M$, the number of subsampled opponents for each candidate, for the league-style algorithm with $N=16$.
@@
Figure 15: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$ that are estimated with the empirical results for the league-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{\Delta}>0$, $\widehat{\Delta}\leq 0$, $\widehat{p}_{\text{cs}}=0$ or $\widehat{p}_{\text{cs}}=1$. To the right of each plot is a histogram for $\widehat{\Delta}$. Each problem is represented by a circle if it is solved correctly by the league-style algorithm with $N=16$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{cs}}=0$ or 11, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{\Delta}$ for such problems.
@@
Figure 16: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{comp}}$ from the knockout-style algorithm and $\widehat{\Delta}$ from the league-style algorithm (both with $N=16$) using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. The following abbreviations are used for the legend: K — knockout, L — league, R — right, W — wrong. For example, “KW-LR (# = 66)” means that there are $6$ problems for which the knockout-style algorithm did wrong while the league-style algorithm did right.